# CardioMOD-Net: A Modal Decomposition–Neural Network Framework for Diagnosis and Prognosis of HFpEF from Echocardiography Cine Loops

### v1

@@ 1 Introduction
Heart failure with preserved ejection fraction (HFpEF) is a major and expanding global health problem, now accounting for more than half of all heart-failure cases and rising steadily in ageing populations [borlaug2020evaluation]. Its clinical complexity derives from the interplay of multiple comorbidities, such as obesity, systemic arterial hypertension, or chronic hyperglycaemia, which frequently coexist and shape distinct patterns of myocardial remodelling. As a consequence, HFpEF does not represent a single disease entity but a spectrum of comorbidity-driven phenotypes that evolve through prolonged subclinical stages before culminating in symptomatic congestion. This heterogeneity complicates both experimental modelling and clinical decision-making and remains a major barrier to identifying early disease or anticipating the transition to symptomatic congestion [borlaug2020evaluation] [akerman2025external]. Diagnosis relies primarily on echocardiography, integrating morphological, functional, and Doppler parameters. However, the early stages of HFpEF often present with subtle or overlapping phenotypes, particularly across comorbidity-driven subtypes such as obesity, systemic arterial hypertension, or chronic hyperglycaemia. This complexity reduces the sensitivity of conventional diagnostic approaches and frequently results in indeterminate assessments, even when established diagnostic scores are applied to standard echocardiographic images [cassianni2024automated] [akerman2023automated].
An even greater unmet need lies in prognosis. Clinicians currently lack tools capable of estimating when a patient with risk factors or early HFpEF will progress to decompensation, and even advanced diagnostic algorithms offer limited temporal prognostic precision [akerman2025external] [cassianni2024automated]. Current prognostic models typically provide risk categories rather than specific timelines, and their performance is hindered by the scarcity of longitudinal imaging data.
Artificial Intelligence (AI) offers the potential to capture subtle features of cardiac motion that are not accessible to visual assessment. Yet most echocardiography-based AI studies focus on narrow diagnostic tasks, rely on very large datasets or pre-measured echocardiographic parameters, and treat diagnosis and prognosis as separate problems rather than components of a unified workflow [cassianni2024automated] [akerman2023automated] [kobayashi2022machine]. There is currently no unified framework capable of performing multiclass cardiac phenotyping and continuous prediction of HFpEF onset from standard echocardiographic cine loops.
@@ 2.1 Data acquisition and experimental groups
Echocardiography videos were obtained from 37 C57BL/6 mice across four experimental groups: control (CTL), hyperglycaemic (HG), obesity (OB), and systemic arterial hypertension (SAH), as we previously described [villalba2025unraveling]. All procedures followed approved institutional animal care protocols (PROEX177/17) and ARRIVE guidelines. Standard two-dimensional parasternal long-axis (PLAX) cine loops were acquired using a high-frequency ultrasound system (Vevo 2100, 40-MHz probe) [villalba2022non]. Videos had a mean resolution of $692\times 599$ pixels and contained on average $284\pm 44$ frames per sequence. For each mouse, multiple echocardiography sequences were collected at different ages. HFpEF onset was defined according to established criteria: pulmonary congestion, measure by lung ultrasound, together with echocardiographic signs of diastolic dysfunction in the presence of preserved left ventricular ejection fraction, and recorded as the age (in weeks) at which each mouse first met this threshold [villalba2025unraveling].
@@ 2.3 AI framework and feature extraction
The unified tool, CardioMOD-Net, integrates two previously published systems for diagnosis and prognosis into a single architecture [groun2025eigenhearts] [bell2025automatic]. Each cine loop was processed using Higher Order Dynamic Mode Decomposition (HODMD) to extract spatiotemporal modes that encode cardiac motion patterns [groun2022higher]. These HODMD-derived features were used to build a shared latent representation supporting both diagnostic and prognostic tasks.
@@ 2.4 Diagnosis module
Heart-state classification (CTL, HG, OB, SAH) was performed using a lightweight Vision Transformer incorporating Masked Autoencoder (MAE)–based strategies [das2024limited] trained on the HODMD feature representation. This design reduces dependence on large datasets and mitigates overfitting by exploiting temporal coherence within cine loops. Diagnostic performance was evaluated using class-specific accuracy and a four-class confusion matrix.
@@ 4 Discussion
In this study, we present CardioMOD-Net, an AI tool capable of performing multiclass cardiac diagnosis and continuous prediction of HFpEF onset directly from standard PLAX cine loops, addressing methodological gaps that remain unfilled in current echocardiography-based AI. Although recent advances have demonstrated that deep learning can identify HFpEF from single apical four-chamber videos in humans, these approaches remain largely restricted to binary classification and do not provide temporal estimates of disease progression [cassianni2024automated] [akerman2023automated]. Likewise, large multi-task echocardiographic systems trained on millions of clinical images offer comprehensive diagnostic outputs but require extensive datasets and do not address comorbidity-specific phenotyping or time-to-event prediction [akerman2025external]. Our findings show that both diagnosis and prognosis can be extracted from a unified feature representation even under small-data conditions, supporting the feasibility of efficient, task-integrated AI models.
@@
Figure
Figure 1: Diagnosis and prognosis results obtained with the CardioMOD-Net tool (Cardiovascular Modal Decomposition - Neural Network). A, Schematic depicting the workflow used by this AI-based tool for heart disease recognition. Top, training of the tool for diagnosis and prognosis; bottom, diagnosis and prognosis performed by the tool after it is trained. B, Summary of the main characteristics of echocardiography images in our database. C, Diagnosis results obtained with the tool using the test sequences. D, Confusion matrix obtained with the developed tool using test sequences broken down into age intervals. Top, global confusion matrix; bottom, confusion matrices by age intervals. E, Prognosis results obtained with the tool using the test sequences. CTL, control; HG, hyperglycaemic; OB, obesity; SAH, systemic arterial hypertension; RMSE, Root Mean Squared Error.

### v2

@@ 1 Introduction
Heart failure with preserved ejection fraction (HFpEF) is a major and expanding global health problem, now accounting for more than half of all heart-failure cases and rising steadily in ageing populations [4]. Its clinical complexity derives from the interplay of multiple comorbidities, such as obesity, systemic arterial hypertension, or chronic hyperglycaemia, which frequently coexist and shape distinct patterns of myocardial remodelling. As a consequence, HFpEF does not represent a single disease entity but a spectrum of comorbidity-driven phenotypes that evolve through prolonged subclinical stages before culminating in symptomatic congestion. This heterogeneity complicates both experimental modelling and clinical decision-making and remains a major barrier to identifying early disease or anticipating the transition to symptomatic congestion [4] [1]. Diagnosis relies primarily on echocardiography, integrating morphological, functional, and Doppler parameters. However, the early stages of HFpEF often present with subtle or overlapping phenotypes, particularly across comorbidity-driven subtypes such as obesity, systemic arterial hypertension, or chronic hyperglycaemia. This complexity reduces the sensitivity of conventional diagnostic approaches and frequently results in indeterminate assessments, even when established diagnostic scores are applied to standard echocardiographic images [5] [2].
An even greater unmet need lies in prognosis. Clinicians currently lack tools capable of estimating when a patient with risk factors or early HFpEF will progress to decompensation, and even advanced diagnostic algorithms offer limited temporal prognostic precision [1] [5]. Current prognostic models typically provide risk categories rather than specific timelines, and their performance is hindered by the scarcity of longitudinal imaging data.
Artificial Intelligence (AI) offers the potential to capture subtle features of cardiac motion that are not accessible to visual assessment. Yet most echocardiography-based AI studies focus on narrow diagnostic tasks, rely on very large datasets or pre-measured echocardiographic parameters, and treat diagnosis and prognosis as separate problems rather than components of a unified workflow [5] [2] [9]. There is currently no unified framework capable of performing multiclass cardiac phenotyping and continuous prediction of HFpEF onset from standard echocardiographic cine loops.
@@ 2.1 Data acquisition and experimental groups
Echocardiography videos were obtained from 37 C57BL/6 mice across four experimental groups: control (CTL), hyperglycaemic (HG), obesity (OB), and systemic arterial hypertension (SAH), as we previously described [11]. All procedures followed approved institutional animal care protocols (PROEX177/17) and ARRIVE guidelines. Standard two-dimensional parasternal long-axis (PLAX) cine loops were acquired using a high-frequency ultrasound system (Vevo 2100, 40-MHz probe) [10]. Videos had a mean resolution of $692\times 599$ pixels and contained on average $284\pm 44$ frames per sequence. For each mouse, multiple echocardiography sequences were collected at different ages. HFpEF onset was defined according to established criteria: pulmonary congestion, measure by lung ultrasound, together with echocardiographic signs of diastolic dysfunction in the presence of preserved left ventricular ejection fraction, and recorded as the age (in weeks) at which each mouse first met this threshold [11].
@@ 2.3 AI framework and feature extraction
The unified tool, CardioMOD-Net, integrates two previously published systems for diagnosis and prognosis into a single architecture [7] [3]. Each cine loop was processed using Higher Order Dynamic Mode Decomposition (HODMD) to extract spatiotemporal modes that encode cardiac motion patterns [8]. These HODMD-derived features were used to build a shared latent representation supporting both diagnostic and prognostic tasks.
@@ 2.4 Diagnosis module
Heart-state classification (CTL, HG, OB, SAH) was performed using a lightweight Vision Transformer incorporating Masked Autoencoder (MAE)–based strategies [6] trained on the HODMD feature representation. This design reduces dependence on large datasets and mitigates overfitting by exploiting temporal coherence within cine loops. Diagnostic performance was evaluated using class-specific accuracy and a four-class confusion matrix.
@@ 4 Discussion
In this study, we present CardioMOD-Net, an AI tool capable of performing multiclass cardiac diagnosis and continuous prediction of HFpEF onset directly from standard PLAX cine loops, addressing methodological gaps that remain unfilled in current echocardiography-based AI. Although recent advances have demonstrated that deep learning can identify HFpEF from single apical four-chamber videos in humans, these approaches remain largely restricted to binary classification and do not provide temporal estimates of disease progression [5] [2]. Likewise, large multi-task echocardiographic systems trained on millions of clinical images offer comprehensive diagnostic outputs but require extensive datasets and do not address comorbidity-specific phenotyping or time-to-event prediction [1]. Our findings show that both diagnosis and prognosis can be extracted from a unified feature representation even under small-data conditions, supporting the feasibility of efficient, task-integrated AI models.